# SlideSLAM: Sparse, Lightweight, Decentralized Metric-Semantic SLAM for Multi-Robot Navigation

### v4

@@ Abstract
This paper develops a real-time decentralized metric-semantic Simultaneous Localization and Mapping (SLAM) algorithm framework that enables a heterogeneous robot team to collaboratively construct object-based metric-semantic maps of 3D environments featuring indoor, urban, and forests without relying on GPS. The framework integrates a data-driven front-end for instance segmentation from either RGBD cameras or LiDARs and a custom back-end for optimizing robot trajectories and object landmarks in the map. To allow multiple robots to merge their information, we design semantics-driven place recognition algorithms that leverage the informativeness and viewpoint invariance of the object-level metric-semantic map for inter-robot loop closure detection. A communication module is designed to track each robot’s observations and those of other robots whenever communication links are available. Our framework enables real-time decentralized operations onboard robots, allowing them to opportunistically leverage communication. We integrate the proposed framework with the autonomous navigation and exploration systems of three types of aerial and ground robots, conducting extensive experiments in a variety of indoor and outdoor environments. These experiments demonstrate the framework’s accuracy in inter-robot localization and object mapping, along with its moderate demands on computation, storage, and communication resources. The framework is open-sourced and available as a modular stack for object-level metric-semantic SLAM, suitable for both single-agent and multi-robot scenarios. The project website and code can be found at https://xurobotics.github.io/slideslam/ and https://github.com/XuRobotics/SLIDE_SLAM, respectively.
@@ I Introduction
We develop a multi-robot decentralized metric-semantic SLAM system for autonomous navigation and exploration. We model the environment using hierarchical metric-semantic maps containing three levels of abstraction: high-level, mid-level, and low-level. The high-level abstraction stores a sparse landmark map that explicitly models objects and can be used for task planning and long-time-horizon SLAM. The mid-level abstraction maintains an optional accumulated segmented point cloud map that can be activated to provide detailed semantic information on demand. The low-level abstraction is a volumetric map over short durations, which is used for navigation and obstacle avoidance.
@@
Algorithm: We develop a real-time decentralized metric-semantic SLAM framework that supports heterogeneous aerial and ground robots. The framework features: (a) a flexible front-end that supports LiDAR and RGBD sensors, and can perform both closed-set as well as open-vocabulary [^1] In this paper, we consider the term open-vocabulary (object detection or semantic mapping) to encompass what is also referred to as open-world and open-set (object detection or semantic mapping), to ensure consistency throughout the paper. object detection, (b) an efficient back-end customized for our object-level metric-semantic map representation, (c) semantics-driven place recognition algorithms that utilize sparse and lightweight object-level semantic maps for map merging, and (d) a decentralized multi-robot collaboration module that facilitates information sharing, even under intermittent communication conditions.
@@
System integration: We integrate and deploy the proposed framework on a heterogeneous team of robots, demonstrating its capacity to enable semantics-in-the-loop autonomous navigation and exploration in various indoor and outdoor environments. The system operates in real time onboard SWaP-constrained robots, utilizing limited computation and memory resources.
Experiments: We conduct extensive real-world experiments and provide thorough empirical results and analysis that highlight the efficiency, accuracy, and robustness of our system.
@@
Figure 3: System Diagram. Our system takes in data streams from each robot’s onboard sensors, which can be either an RGBD camera or a LiDAR, and performs instance segmentation to extract semantic object features. Meanwhile, low-level odometry, either Visual-Inertial Odometry (VIO) [17] or LiDAR-Inertial Odometry (LIO) [18], provides relative-motion estimates between consecutive key poses. Next, the metric-semantic SLAM algorithm takes in such semantic observations and relative motion estimates, and constructs a factor graph consisting of both robot pose nodes and semantic object landmark nodes. Meanwhile, our multi-robot communication module (see Fig. 4) opportunistically leverages communication to share lightweight semantic observations among robots in a decentralized way. Based on this shared information, our metric-semantic place recognition algorithm constantly checks for possible inter-robot loop closures at a fixed rate. Once a loop closure is detected, the resulting transformation between each pair of robots is used to transform all observations into each robot’s reference frame. These observations are then added to their own factor graphs, forming a merged metric-semantic map. The output of the metric-semantic SLAM module will be taken as input for the planning module, which will plan both high-level exploration and low-level dynamically feasible trajectories that guide the robot to explore and navigate the environment. Note that the entire perception-action loop runs in a decentralized manner onboard each robot. Besides the obvious differences in control algorithms, the planning modules and the front-end processing algorithms are also different across each robot platform. This is due to the need to accommodate the differences in sensing modalities (RGBD and LiDAR), operating environments (indoor, urban, and forests), and traversal modes (ground and aerial). However, the core metric-semantic SLAM framework remains the same.
@@ II-A Metric-semantic SLAM
Unlike traditional SLAM, metric-semantic SLAM constructs a map that encodes geometric features and semantic information on objects of interest. Metric-semantic SLAM has gained significant success and popularity in the past decade [19], driven by the rapid development of deep learning techniques that extract semantic information from sensor data. A variety of map representations are used in existing metric-semantic SLAM literature. Some use dense semantic maps, such as meshes [20], volumetric maps [21, 22], surfels [23], and 2.5D grid maps [24, 25]. Others use sparse object-level maps with prior information of object shapes, such as centroids [11], cubes [8], ellipsoids [9], cylinders [26], structured object models with prior shape constraints [27, 7], and mesh-based object models [28, 12]. While dense metric-semantic maps are suitable for obstacle avoidance, their computational and memory demands, especially in SWaP-constrained platforms, can be significant. Thus, sparse object-level maps are desirable for real-time downstream tasks such as active information gathering, object manipulation and multi-robot collaboration. This is because semantic information is essential for robots to perceive their environment intelligently and execute tasks accordingly. Sparsity helps robots reduce resource demands on computation, storage, and communication. Therefore, our work seeks to develop a metric-semantic SLAM framework that utilizes sparse explicitly modeled objects and supports real-time decentralized operations within a heterogeneous robot team.
@@ II-D Semantics-in-the-loop navigation and exploration
Most autonomous navigation systems rely on geometric maps. Although such maps are reliable for obstacle avoidance, they lack semantic information that robots can leverage to improve their state estimation and planning while navigating the environment. Object-based landmarks help minimize odometry drifts, thus benefiting autonomous navigation if executed onboard in real-time [6]. Recent work goes beyond improving odometry but uses semantic information for active uncertainty reduction planning [48, 49, 3], and multi-robot collaborative planning [25]. The use of semantic information in the active SLAM is limited [50]. Some representative work in this field includes [51, 52, 1]. However, there has not yet been a real-time decentralized metric-semantic SLAM framework that enables a team of heterogeneous robots to explore GPS-denied environments while reasoning about the semantic information and leveraging it to facilitate multi-robot collaboration. Our work seeks to develop such a framework.
@@ IV-B Map representation
Table II: Requirements for map representation. Generic implies that the map must be adaptable to various sensing modalities and environments. Informative means that the map should contain both metric and rich semantic information. Sparse indicates that the map needs to be memory efficient for storage and sharing.
@@ IV-E Place recognition and loop closure
The two example algorithms exhibit different behaviors and offer complementary performance across various scenarios. Generally, we recommend using SlideGraph as the primary algorithm, with SlideMatch serving as an alternative option, as further discussed in detail in Section VI.
@@ IV-E1 SlideMatch
The SlideMatch algorithm takes a pair of object-level metric-semantic maps from either the same robot (i.e., historical map and current map for loop closure) or a pair of different robots. It checks within a search region if a valid loop closure is found and outputs the relative transformation between the reference frames of these two maps. This search region encompasses continuous ranges of X and Y positions and yaw angles, where the relative transformation between the two maps may exist. The search procedure is carried out over the search region using a user-defined search resolution for discretization.
@@ IV-E3 Least square optimization
Figure 6: Multi-robot metric-semantic SLAM in an outdoor experiment. Results from an experiment involving three Falcon 4 UAV flights. The upper-right panel provides an orthographic view. The trajectories, marked by red, blue, and yellow bold lines, start from different positions (red, blue, and yellow dots) but converge back to the same end location (red dot). Our semantics-driven place recognition algorithm was able to detect inter-robot loop closures, which are then used to merge the maps. In the map, vehicle models are represented by blue cuboids, and models for tree trunks or light poles are shown by magenta cylinders.
@@ IV-F Decentralized graph optimization for multi-robot SLAM
One natural question that may arise is why do we share such lightweight observations between robots instead of sharing marginalized factor graphs (e.g., as in [13, 14])? We justify our design choice by the fact that the object-level metric-semantic map that we use is significantly sparser than traditional geometric representations such as point, line, or planar features. In addition, the number of objects detected at each pose is much smaller compared to geometric features (e.g., corner points). In terms of performance, our representation allows us to be memory and computationally efficient, as supported by the experimental results in Table VII. In terms of statistical consistency, our approach does not suffer from spurious information loss due to linearization.
@@ V-A Robot platform overview
! Sensor Environment Platform Autonomy Unknown TF Num. Robots Experiment 1 RGBD+LiDAR Indoor+Outdoor Aerial+Ground No Yes 7 Sequential Experiment 2 RGBD Indoor Aerial+Ground Yes Yes 4 Simultaneous Experiment 3 LiDAR Outdoor Aerial Yes Yes 3 Sequential Experiment 4 LiDAR Forest Aerial No Yes 3 Sequential
Table IV: Experiment design. The main purpose of these experiments is to demonstrate the versatility of our system and gauge performance in real-world applications across a heterogeneous set of robot platforms. Each set is tailored to test different aspects of the system.
@@ V-C Experiment design
We designed and conducted a set of experiments to evaluate our system, along with its individual modules. To align with the system design attributes outlined in Section IV-B, we organized four distinct sets of experiments. A summary of these experiments is provided in Table IV. Here, we provide an overview of each experiment:
@@ V-C3 Experiment 3
A single-robot outdoor autonomous experiment where the Falcon 4 operates fully autonomously to map an outdoor region. The purpose of this experiment is to demonstrate that our system can operate efficiently when paired with autonomy stack on resourced-constrained UAVs in large-scale outdoor environments.
@@ V-C4 Experiment 4
The experiments in the urban environment were done at the University of Pennsylvania Pennovation campus, located in the U.S. city of Philadelphia, with the Falcon 4 mainly mapping cars, tree trunks, and light poles across different parking lots and the Falcon 250 and Scarabs mapping chairs, tables, monitors, and televisions inside the lab buildings. Forest experiments were conducted in the Wharton State Forest with the Falcon 4 platform mapping tree trunks.
V-D Evaluation metrics and section outline
The following metrics are chosen to evaluate the performance of our system: (1) Inter-robot localization accuracy. (2) Precision, recall, and F-1 score for object mapping. (3) Runtime and communication bandwidth requirements.
The rest of the section is structured as follows: It begins with the demonstration of indoor and outdoor mapping without inter-robot loop closures (experiment 1.a), as detailed in Section V-E. Next, in Section V-F, we present and analyze results on experiments conducted in urban and forest environments (experiment 1.a and experiment 4), where we demonstrate inter-robot loop closure and map merging. The discussion then extends to experiments involving autonomous aerial and ground robots in indoor settings (experiment 2), as detailed Section V-G. This is followed by Section V-H, which presents results on cross-sensing-modality experiment using robots equipped with LiDAR and RGBD sensors (experiment 1.b). Additionally, we report our results on object mapping accuracy in Section V-I. Finally, we present an analysis of communication and runtime metrics in Section V-K.
! Experiment (Sensor) Position Err. (SM) [m] Yaw Err. (SM) [] Position Err. (SG) [m] Yaw Err. (SG) [] Outdoor-Urban Falcon 4 #0 to Falcon 4 #1 (LiDAR) 0.235 -0.2 0.074 0.3 Outdoor-Urban Falcon 4 #1 to Falcon 4 #2 (LiDAR) 0.159 -0.3 0.240 0.9 Outdoor-Urban Falcon 4 #0 to Falcon 4 #2 (LiDAR) 0.135 -0.9 0.848 0.7 Outdoor-Forest Falcon 4 #0 to Falcon 4 #2 (LiDAR) 0.040 3.1 0.101 0.3 Outdoor-Forest Falcon 4 #1 to Falcon 4 #2 (LiDAR) 0.159 3.2 0.049 -3.0 Outdoor-Forest Falcon 4 #0 to Falcon 4 #1 (LiDAR) 0.054 -0.8 0.126 -2.6 Cross-sensing-modality (LiDAR-RGBD) 0.542 -4.0 0.578 2.0 Indoor Scarab 45 to Scarab 40 (RGBD) 0.221 -3.4 0.952 -2.3 Indoor Scarab 40 to Scarab 41 (RGBD) 0.165 0.0 0.253 0.1 Indoor Scarab 45 to Scarab 41 (RGBD) 0.460 3.7 1.040 -6.1 Indoor aerial-ground (RGBD) 0.245 -2.4
Table V: Inter-robot localization errors. SM and SG refer to the SlideMatch and SlideGraph algorithms, respectively. Position error is calculated as the L2 norm of the errors in the X, Y, and Z coordinates. In general, the SlideGraph algorithm is more efficient and requires less parameter tuning compared to SlideMatch, while still providing comparable results. However, both algorithms encounter difficulties in scenarios where object landmarks are noisy, particularly with RGBD-based sensing. SlideGraph is relatively more prone to failures in such conditions due to insufficient pairwise consistency among object landmarks.
@@ V-E Indoor-outdoor LiDAR-RGBD metric-semantic SLAM without inter-robot loop closures
Fig. 2 shows qualitative results on metric-semantic SLAM for experiment 1.a as described in Section V-C. Outdoor objects such as cars, tree trunks, and light poles and indoor objects such as chairs, tables, and monitors are detected and mapped using our sparse metric-semantic map representation. Object models for vehicles are represented by blue cuboids, while tree trunks and light pole landmarks are shown as pink cylinders. Chairs, tables, and monitors are initially mapped using an ellipsoid model and subsequently replaced with a Computer-Aided Design (CAD) model. These models are colored blue, green, and red, respectively, according to their detected semantic class. In this experiment, we showcase the capabilities of our metric-semantic SLAM framework in integrating data from multiple robots equipped with either LiDARs or RGBD cameras, enabling the construction of a unified metric-semantic map that includes both indoor and outdoor areas.

### v5

@@ Abstract
[…]
Index Terms:
Metric-Semantic SLAM; Multi-Robot Systems; Place Recognition; Aerial Systems; Heterogeneous Robots
@@ I Introduction
We develop a multi-robot decentralized metric-semantic SLAM system that is open-sourced [^1] Code: https://github.com/KumarRobotics/SLIDE_SLAM. Project website: https://xurobotics.github.io/slideslam/. and suitable for real-time autonomous navigation and exploration with resource-constrained heterogeneous robots. We model the environment using hierarchical metric-semantic maps containing three levels of abstraction: high-level, mid-level, and low-level. The high-level abstraction stores a sparse landmark map that explicitly models objects and can be used for task planning and long-time-horizon SLAM. The mid-level abstraction maintains an optional accumulated segmented point cloud map that can be activated to provide detailed semantic information on demand. The low-level abstraction is a volumetric map over short durations, which is used for navigation and obstacle avoidance.
@@
Algorithm: We develop a real-time decentralized metric-semantic SLAM framework that supports heterogeneous aerial and ground robots. The framework features:
  1. (a)
    a flexible front-end that supports LiDAR and RGBD sensors, and can perform both closed-set as well as open-vocabulary object detection,
  2. (b)
    an efficient back-end customized for our object-level metric-semantic map representation,
  3. (c)
    semantics-driven place recognition algorithms that utilize sparse and lightweight object-level semantic maps for map merging,
  4. (d)
    a decentralized multi-robot collaboration module that facilitates information sharing, even under intermittent communication conditions.
@@
System integration: We integrate and deploy the proposed framework on a heterogeneous team of robots, demonstrating its capacity to enable semantics-in-the-loop autonomous navigation and exploration in various indoor and outdoor environments. The system operates in real time onboard SWaP-constrained robots, while maintaining moderate computation and memory demands.
[…]
@@
Figure 3: System Diagram. Our system takes in data streams from each robot’s onboard sensors, which can be either an RGBD camera or a LiDAR, and performs instance segmentation to extract semantic object features. Meanwhile, low-level odometry, either Visual-Inertial Odometry (VIO) [17] or LiDAR-Inertial Odometry (LIO) [18], provides relative-motion estimates between consecutive key poses. Next, the metric-semantic SLAM algorithm takes in such semantic observations and relative motion estimates, and constructs a factor graph consisting of both robot pose nodes and semantic object landmark nodes. Meanwhile, our multi-robot communication module (see Fig. 4) opportunistically leverages communication to share lightweight semantic observations among robots in a decentralized way. Based on this shared information, our metric-semantic place recognition algorithm constantly checks for possible inter-robot loop closures at a fixed rate. Once a loop closure is detected, the resulting transformation between each pair of robots is used to transform all observations into each robot’s reference frame. These observations are then added to their own factor graphs, forming a merged metric-semantic map. Note that the entire perception-action loop runs in a decentralized manner onboard each robot. Besides the obvious differences in control algorithms, the planning modules and the front-end processing algorithms are also different across each robot platform. This is due to the need to accommodate the differences in sensing modalities (RGBD and LiDAR), operating environments (indoor, urban, and forests), and traversal modes (ground and aerial). However, the core metric-semantic SLAM framework remains the same.
@@ II-A Metric-semantic SLAM
Unlike traditional SLAM, metric-semantic SLAM constructs a map that encodes geometric features and semantic information on objects of interest. Metric-semantic SLAM has gained significant success and popularity in the past decade [19], driven by the rapid development of deep learning techniques that extract semantic information from sensor data. A variety of map representations are used in existing metric-semantic SLAM literature. Some use dense semantic maps, such as meshes [20], volumetric maps [21, 22], surfels [23], and 2.5D grid maps [24, 25, 26]. Others use sparse object-level maps with prior information of object shapes, such as centroids [11], cubes [8], ellipsoids [9], cylinders [27], structured object models with prior shape constraints [28, 7], and mesh-based object models [29, 12]. While dense metric-semantic maps are suitable for obstacle avoidance, their computational and memory demands, especially in SWaP-constrained platforms, can be significant. Thus, sparse object-level maps are desirable for real-time downstream tasks such as active information gathering, object manipulation and multi-robot collaboration. This is because semantic information is essential for robots to perceive their environment intelligently and execute tasks accordingly. Sparsity helps robots reduce resource demands on computation, storage, and communication. Therefore, our work seeks to develop a metric-semantic SLAM framework that utilizes sparse explicitly modeled objects and supports real-time decentralized operations within a heterogeneous robot team.
@@ II-D Semantics-in-the-loop navigation and exploration
Most autonomous navigation systems rely on geometric maps. Although such maps are reliable for obstacle avoidance, they lack semantic information that robots can leverage to improve their state estimation and planning while navigating the environment. Object-based landmarks help minimize odometry drifts, thus benefiting autonomous navigation if executed onboard in real-time [6]. Recent work goes beyond improving odometry but uses semantic information for active uncertainty reduction planning [49, 50, 3], and multi-robot collaborative planning [26]. The use of semantic information in the active SLAM is relatively uncommon [51]. Some representative work in this field includes [52, 53]. However, there has not yet been a real-time decentralized metric-semantic SLAM framework that enables a team of heterogeneous robots to explore GPS-denied environments while reasoning about the semantic information and leveraging it to facilitate multi-robot collaboration. Our work seeks to develop such a framework.
@@ IV-B Map representation
Table II: Requirements for map representation. Generic implies that the map must be adaptable to various sensing modalities and environments. Informative means that the map should contain both metric and semantic information. Sparse indicates that the map needs to be memory efficient for storage and sharing.
@@ IV-E Place recognition and loop closure
The two example algorithms exhibit different behaviors and offer complementary performance across various scenarios. Generally, we recommend using SlideGraph as the primary algorithm, with SlideMatch serving as an alternative option, as further discussed in detail in Section V-L.
@@ IV-E1 SlideMatch
The SlideMatch algorithm takes a pair of object-level metric-semantic maps from either the same robot (i.e., historical map and current map for loop closure) or a pair of different robots. Similar to our previous work in [63], SlideMatch algorithm checks within a search region if a valid loop closure is found and outputs the relative transformation between the reference frames of these two maps. Unlike our previous work [63], here we also leverage additional information such as semantic labels and dimensions of objects in the maps. The search region of the algorithm encompasses continuous ranges of X and Y positions and yaw angles, where the relative transformation between the two maps may exist. The search procedure is carried out over the search region using a user-defined search resolution for discretization.
@@ IV-E3 Least square optimization
Figure 6: Multi-robot metric-semantic SLAM in an outdoor experiment. Results from an experiment involving three Falcon 4 UAV flights. The trajectories, marked by red, blue, and yellow bold lines, start from different positions (red, blue, and yellow dots) but converge back to the same end location (red dot). Our semantics-driven place recognition algorithm was able to detect inter-robot loop closures, which are then used to merge the maps. In the map, vehicle models are represented by blue cuboids, and models for tree trunks or light poles are shown by magenta cylinders.
@@ IV-F Decentralized graph optimization for multi-robot SLAM
One natural question that may arise is why do we share such lightweight observations between robots instead of sharing marginalized factor graphs (e.g., as in [13, 14])? We justify our design choice by the fact that the object-level metric-semantic map that we use is significantly sparser than traditional geometric representations such as point, line, or planar features. In addition, the number of objects detected at each pose is much smaller compared to geometric features (e.g., corner points). In terms of performance, our representation allows us to be memory and computationally efficient, as supported by the experimental results in Table VI. In terms of statistical consistency, our approach does not suffer from spurious information loss due to linearization.
@@ V-C Experiment design
We designed and conducted a set of experiments to evaluate our framework and its modules. We conducted four sets of experiments. Here, we provide an overview of each experiment:
@@ V-C4 Experiment 4
In the final experiment, we tested the proposed framework on the publicly available Semantic KITTI dataset. This allowed us to demonstrate the algorithm’s versatility and effectiveness in scenarios beyond our custom robot platforms and experimental setups.
The urban environment experiments were conducted at the University of Pennsylvania’s Pennovation campus in Pennsylvania, USA. Forest experiments were conducted in the Wharton State Forest in New Jersey, USA.
V-D Evaluation metrics
The following metrics are chosen to evaluate the performance of our system: (1) Inter-robot localization accuracy. (2) Precision, recall, and F-1 score for object mapping. (3) Runtime and communication bandwidth requirements. In the rest of this section, we present a detailed analysis of the results for each experiment.
! Experiment (Sensor) Position Err. (SM) [m] Yaw Err. (SM) [] Position Err. (SG) [m] Yaw Err. (SG) [] Outdoor-Urban Falcon 4 #0 to Falcon 4 #1 (LiDAR) 0.235 -0.2 0.074 0.3 Outdoor-Urban Falcon 4 #1 to Falcon 4 #2 (LiDAR) 0.159 -0.3 0.240 0.9 Outdoor-Urban Falcon 4 #0 to Falcon 4 #2 (LiDAR) 0.135 -0.9 0.848 0.7 Outdoor-Forest Falcon 4 #0 to Falcon 4 #2 (LiDAR) 0.040 3.1 0.101 0.3 Outdoor-Forest Falcon 4 #1 to Falcon 4 #2 (LiDAR) 0.159 3.2 0.049 -3.0 Outdoor-Forest Falcon 4 #0 to Falcon 4 #1 (LiDAR) 0.054 -0.8 0.126 -2.6 Cross Sensing Modality (LiDAR-RGBD) 0.542 -4.0 0.578 2.0 Indoor Scarab 45 to Scarab 40 (RGBD) 0.221 -3.4 0.952 -2.3 Indoor Scarab 40 to Scarab 41 (RGBD) 0.165 0.0 0.253 0.1 Indoor Scarab 45 to Scarab 41 (RGBD) 0.460 3.7 1.040 -6.1 Indoor Aerial-Ground (RGBD) 0.245 -2.4
Table IV: Inter-robot localization errors. SM and SG refer to the SlideMatch and SlideGraph algorithms, respectively. Position error is calculated as the L2 norm of the errors in the X, Y, and Z coordinates of the relative transformation between two robots’ reference frames. In general, the SlideGraph algorithm is more efficient and requires less parameter tuning compared to SlideMatch, while still providing comparable results. However, both algorithms encounter difficulties in scenarios where object landmarks are noisy, particularly with RGBD-based sensing. SlideGraph is relatively more prone to failures in such conditions due to insufficient pairwise consistency among object landmarks.
@@ V-E Indoor-outdoor LiDAR-RGBD metric-semantic SLAM without inter-robot loop closures
Fig. 2 shows qualitative results on metric-semantic SLAM for experiment 1.a as described in Section V-C. Outdoor objects such as cars, tree trunks, and light poles and indoor objects such as chairs, tables, and monitors are detected and mapped using our sparse metric-semantic map representation. Object models for vehicles are represented by blue cuboids, while tree trunks and light pole landmarks are shown as pink cylinders. Chairs, tables, and monitors are initially mapped using an ellipsoid model and subsequently replaced with a Computer-Aided Design (CAD) model. These models are colored blue, green, and red, respectively, according to their detected semantic class. This experiment shows the capabilities of our metric-semantic SLAM framework in integrating data from multiple robots equipped with either LiDARs or RGBD cameras, enabling the construction of a unified metric-semantic map that includes both indoor and outdoor areas.